The length and width of a rectangle are in the ratio 5:4. The perimeter of the rectangle is 126 inches. What are the length and width of the rectangle.

Answer by stanbon(75887)

The length and width of a rectangle are in the ratio 5:4. The perimeter of the rectangle is 126 inches. What are the length and width of the rectangle.
----------------
5:4 is the same as 5x to 4x where 5x is the length and 4x is the width.
----------
P = 2(L + W)
126 = 2(5x+4x)
63 = 9x
x = 7
-----
Length = 5x = 5*7 = 35 in
Width = 4x = 4*7 = 28 in
-----------------------
Cheers,
Stan H.